# Convex Physics Informed Neural Networks for the Monge-Ampère Optimal Transport Problem

### v1

@@ 1. Introduction
In the spirit of Monge [20], the optimal transport problem describes the problem of moving bulk material from a source (supplier) to a target destination location (customer), see, e.g., [7, 22, 23].
From the numerical viewpoint, many numerical methods from classical scientific computing have been advocated in the literature, involving, e.g., finite differences approaches [3, 5, 10] and finite element methods [4, 25, 30]. On the other hand, the emergence of deep neural networks has opened many avenues of investigation to approximate partial differential equations, see, e.g., [13, 16, 18]. In particular, Physics Informed Neural Networks (PINNs) have shown their capabilities to approximate solutions of partial differential equations [26]. Several extensions of PINNs are currently being investigated, see, e.g., consistent PINNs [6] or variational PINNS [17]. Applications of PINNs to the classical Monge-Ampère equation (with Dirichlet boundary conditions) can be found in [21], while a first tentative to solve the optimal transport problem is presented in [28].
We aim here at solving the following optimal transport problem. Let $\mathcal{X}$ and $\mathcal{Y}$ be two bounded, smooth open sets in $\mathbb{R}^{2}$, equipped respectively with probability measures $\mu(dx)=f(x)\,dx$ and $\nu(dy)=g(y)\,dy$, where the density functions $f$ and $g$ are supported on $\mathcal{X}$ and $\mathcal{Y}$, respectively. We further assume that $f$ and $g$ are strictly positive and bounded on their respective supports. Brenier’s Theorem, see, e.g., [14], implies that the optimal transport map, with respect to a quadratic cost, is given by the gradient of a convex potential $u$, which satisfies the generalized Monge-Ampère equation: find $u:\mathcal{X}\rightarrow\mathbb{R}$ such that
@@ 2. Mathematical Model
$\Omega$$x$$\mathcal{X}$$\partial\mathcal{X}$$\mathcal{Y}$$\partial\mathcal{Y}$$\mathbf{m}(x)$
@@
[…]
@@ 4.1. Implementation details
For the implementation of the PINN solver, the code has been developed in pytorch and the training is performed using input convex neural networks (ICNNs). The neural networks used in the test cases consist of 4 hidden layers with 10 neurons per layer. The activation function employed is the softplus function, defined as $\sigma(x)=\log(1+e^{x})$, chosen because it is an increasing convex function. To guarantee the positivity of the weight matrices $W^{(l)}\in\mathbb{R}^{N_{l+1}\times N_{l}}$, the weights are initialized by squaring them element-wise, i.e., $w_{ij}\leftarrow w_{ij}^{2}$, as suggested in [21]. Note that the alternative approach of clipping the weights, such that $w_{ij}\geq 0$ for all $i,j$, following [2], does not provide satisfactory results in our experiments.
For the optimization method, we use the L-BFGS algorithm. However, given the sensitivity of L-BFGS to initialization, we first perform several iterations of the Adam optimizer to provide a more robust starting point. This approach allows for more stable convergence of the L-BFGS optimizer.
Moreover, ICNNs are notoriously difficult to train [8, 27] and a meaningful initialization is required. In particular, the network parameters are initialized by training the network to approximate the identity map for the gradient of the solution, $\nabla u\approx\text{Id}$, inspired by the work in [28]. The neural networs are trained for $2500$ L-BFGS epochs.
By default, every test case is run with $N_{c}=800$ collocation points on the interior of the domain and $N_{b}=800$ points on the boundary, choice which will be justified in the numerical experiments in the sequel. The constant $C$ is set to 1, unless specified otherwise.
@@ 4.2. Transport of a constant density from a unit disk domain to an ellipse domain
Figure 2 shows how uniformly sampled points in the unit ball are transported to the ellipse, both with the exact solution and the approximate map. It illustrates an appropriate, uniform, repartition of the sampled points (left), and an accurate transport of those points, whose final position coincides well with their exact values (right). Figure 3 visualizes the corresponding vector field $\nabla u_{NN}$, while Figure 4 details the first and second components of $\nabla u_{NN}$, together with the numerical error component-wise. One can observe that the error on both components is well balanced and of the order $10^{-4}$.
@@
Let us study the sensitivity of those results with respect to the number of epochs and number of collocation points. Figure 5 (top and middle) visualizes the behavior of the $L^{2}$ error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ with respect to these factors respectively. The error has been averaged over a sample of $10$ simulations. In particular, when the number of collocation points increases, e.g., from $N_{c}=50$ to $N_{c}=100$, the first $50$ points in each sample remains the same as those for the smaller sample. It shows that the error decreases when either of these quantities increases, albeit with a large sample variation. Finally, Figure 5 (bottom) visualizes the behavior of the $L^{2}$ error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ with respect to the ratio $N_{b}/N_{c}$ between the number of boundary points and the number of collocation points. It shows that the error is smaller for a ratio equal to one, which confirms the initial choice of $N_{b}=N_{c}$ made earlier.
Figure 5. Transport of a constant density from a unit disk domain to an ellipse domain. Top: $L^{2}$ (test) error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ vs number of epochs. Middle: $L^{2}$ (test) error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ vs number of collocation points. Bottom: $L^{2}$ (test) error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ vs ratio $N_{b}/N_{c}$. (Results averaged on $10$ simulations.)
@@ 4.3. Transport of a constant density from an ellipse domain to a rotated ellipse domain
We run the optimising process with $400$ Adam epochs and $1000$ L-BFGS epochs. Figure 6 shows how uniformly sampled points in the original ellipse $\mathcal{X}$ are transported to the rotated ellipse $\mathcal{Y}$; it illustrates an appropriate, uniform, repartition of the sampled points and an accurate transport of those points, whose final position coincides well with their exact values. Figure 7 visualizes the corresponding vector field $\nabla u_{NN}$. Figure 8 visualizes the approximation error on both components of $\nabla u_{NN}$, and highlights that the error is of the order of $10^{-3}$ pointwise.
@@
Figure 7. Transport of a constant density from an ellipse domain to a rotated ellipse domain Visualization of the transport map $\nabla u_{NN}$.
@@ 4.5. Transport of a Gaussian density in the unit square to another Gaussian density in the unit square
where $c_{0}$ is a normalization constant, $\sigma^{2}=0.25$ and ${x}_{0}=(0.25,0.75)$, and $g({x})$ defined similarly, with $\sigma^{2}=0.25$ and ${x}_{0}=(0.75,0.25)$. The histogram in Figure 11 (left) shows the distribution of points sampled from the source Gaussian distribution (4.1) in the unit square and then transported to the target Gaussian distribution through the map $\nabla u_{NN}$ (right). Even if there is no analytical solution, we observe that the distribution of the transported points very closely resembles a Gaussian distribution.
@@ 5. Conclusions and Perspectives
Since we look for a convex solution of the Monge-Ampère equation, input convex neural networks are suitable to obtain appropriate approximations as they are seeking for solutions in the right subspaces. However, input convex neural networks are notoriously difficult to train.
@@
Perspectives include the further incorporation of several sources and targets, via multimodal Gaussian density functions, for real-life applications, as well as the treatment of obstacles. The investigation of the scaling up of the algorithm in a larger number of dimensions is also of interest.

### v2

@@ 1. Introduction
In the field of fully nonlinear equations, the optimal transport problem has raised numerous fundamental questions over the recent years [1, 24, 30]. In particular, it has been linked with analytical models, such as, e.g., the Monge-Ampère equation [9, 14, 19] or the prescribed Jacobian equation [12]. In the spirit of Monge [20], the optimal transport problem describes the problem of moving bulk material from a source (supplier) to a target destination location (customer) [7, 22, 23].
From the numerical viewpoint, many numerical methods from classical scientific computing have been advocated in the literature, involving, e.g., finite differences approaches [3, 5, 10] and finite element methods [4, 25, 31]. On the other hand, the emergence of deep neural networks has opened many avenues of investigation to approximate partial differential equations [13, 16, 18]. In particular, Physics Informed Neural Networks (PINNs) have shown their capabilities to approximate solutions of partial differential equations [26]. Several extensions of PINNs are currently being investigated, e.g., consistent PINNs [6] or variational PINNS [17]. Applications of PINNs to the classical Monge-Ampère equation (with Dirichlet boundary conditions) can be found in [21], while a first tentative to solve the optimal transport problem is presented in [29].
We aim here at solving the following optimal transport problem. Let $\mathcal{X}$ and $\mathcal{Y}$ be two bounded, smooth open sets in $\mathbb{R}^{2}$, equipped respectively with probability measures $\mu(dx)=f(x)\,dx$ and $\nu(dy)=g(y)\,dy$, where the density functions $f$ and $g$ are supported on $\mathcal{X}$ and $\mathcal{Y}$, respectively. We further assume that $f$ and $g$ are strictly positive and bounded on their respective supports. Brenier’s Theorem [14], implies that the optimal transport map, with respect to a quadratic cost, is given by the gradient of a convex potential $u$, which satisfies the generalized Monge-Ampère equation: find $u:\mathcal{X}\rightarrow\mathbb{R}$ such that
@@ 2. Mathematical Model
$\int_{\mathbb{R}^{n}}\chi(\mathbf{s}(x))f(x)\,dx=\int_{\mathbb{R}^{n}}\chi(\mathbf{s}(x))g(\mathbf{s}(x))\det(\nabla\mathbf{s}(x))\,dx.$
@@ 3.2. Input Convex Neural Networks
In [21], which addresses the approximation of the Dirichlet Monge–Ampère problem, the authors have tested the ICNN architecture up to dimension $4$. They have observed that, in theory, the method is applicable in any dimension. However, in practice, training scales poorly with the space dimension, which poses a significant challenge for optimal transport problems in high dimensions. The authors have also explored an alternative approach to enforcing convexity, based on penalizing the diagonal elements of the Hessian matrix. This method is viable in two dimensions but it does not impose convexity in a structural way. As a result, convergence to a convex solution may only occur asymptotically.
Therefore, following the conclusions in [21], ICNNs remain the best networks to enforce the convexity of the solution, despite being hard to train.
@@ 3.3. Physics Informed Neural Networks
The Hausdorff distance used in the loss function to enforce the transport boundary condition can be computationally expensive. However, the method is general and may be applicable to a variety of situations. Furthermore, in some particular cases, the transport boundary condition admits a simpler formulation, for instance when mapping a square domain onto itself. It can be shown that, in such a case, the optimal transport condition is equivalent to a Neumann condition. Therefore, the boundary terms of the loss function can be simplified and the PINNs solver can be significantly optimized.
@@ 4.1. Implementation details
For the implementation of the PINN solver, the code has been developed in pytorch and it is available on request. The input convex neural networks used in the test cases consist of 4 hidden layers with 10 neurons per layer. The activation function employed is the softplus function, defined as $\sigma(x)=\log(1+e^{x})$, and is chosen because it is an increasing strictly convex function. This choice follows the recommendations provided in [21, 29]. To guarantee the positivity of the weight matrices $W^{(l)}\in\mathbb{R}^{N_{l+1}\times N_{l}}$, the weights are initialized by squaring them element-wise, i.e., $w_{ij}\leftarrow w_{ij}^{2}$, as suggested in [21]. Note that the alternative approach of clipping the weights, such that $w_{ij}\geq 0$ for all $i,j$, following [2], does not provide satisfactory results in our experiments.
Following standard practice in PINNs [27], we perform the initial training with the Adam optimizer before switching to the L‑BFGS optimizer to achieve more stable and accurate convergence.
Moreover, ICNNs are notoriously difficult to train [8, 28] and a meaningful initialization is required. In particular, the network parameters are initialized by training the network to approximate the identity map for the gradient of the solution, $\nabla u\approx\text{Id}$, inspired by the work in [29]. Training details are summarized in Table 1. By default, every test case is run with $N_{c}=800$ collocation points on the interior of the domain and $N_{b}=800$ points on the boundary, choice which will be justified in the numerical experiments in the sequel. The constant $C$ is set to 1, unless specified otherwise.
Table 1. L‑BFGS training configuration.
Main optimizer L‑BFGS
Line search strong Wolfe
Sub-iterations per epoch 20
Learning rate 1
Total epochs 100
Batch size Full-batch
@@ 4.2. Transport of a constant density from a unit disk domain to an ellipse domain
Figure 2 shows how uniformly sampled points in the unit ball are transported to the ellipse, both with the exact solution and the approximate map. It illustrates an appropriate, uniform, repartition of the sampled points (left), and an accurate transport of those points, whose final position coincides well with their exact values (right). Figure 3 visualizes the corresponding vector field $\nabla u_{NN}$, while Figure 4 details the first and second components of $\nabla u_{NN}$, together with the numerical error component-wise. We compare how the L-BFGS performs against the Adam optimizer. In order to do so, we run ten independent trainings, each of them with a different random draw of collocation and boundary points and a different random initialization of network parameters. Figure 5 demonstrates that the L‑BFGS solver allows the loss function to decrease more rapidly but also to reach a lower final error than the Adam solver. Table 2 reports the corresponding averages over these ten runs.
@@
Refer to caption Refer to caption
Figure 5. Transport of a constant density from a unit disk domain to an ellipse domain. Left: Loss function vs epochs. Right: $\|\nabla u-\nabla u_{NN}\|_{L^{2}}$ vs epochs, with the error computed on a validation set. For the L-BFGS optimizer, the total epoch count is taken as the number of (outer) L‑BFGS iterations times the number of inner sub‐iterations. Shaded bands show the 5th–95th percentile range over ten independent runs.
Table 2. Comparison of the Adam vs. L‑BFGS optimizers. Provided values are the average over ten independent runs, (together with standard deviation).
Optimizer Final Loss $L^{2}$ Error (Train) $L^{2}$ Error (Test)
Adam 0.00150(0.0043)0.00150\,(0.0043) 0.02399(0.00327)0.02399\,(0.00327) 0.02504(0.00389)0.02504\,(0.00389)
L‑BFGS $1.13\cdot 10^{-5}\,(1.39\cdot 10^{-5})$ 0.00267(0.00376)0.00267\,(0.00376) 0.00268(0.00376)0.00268\,(0.00376)
[…]
@@ 4.3. Transport of a constant density from an ellipse domain to a rotated ellipse domain
We run the optimising process with $400$ Adam epochs and $1000$ L-BFGS epochs. Figure 7 shows how uniformly sampled points in the original ellipse $\mathcal{X}$ are transported to the rotated ellipse $\mathcal{Y}$; it illustrates an appropriate, uniform, repartition of the sampled points and an accurate transport of those points, whose final position coincides well with their exact values. Figure 8 visualizes the corresponding vector field $\nabla u_{NN}$. Figure 9 visualizes the approximation error on both components of $\nabla u_{NN}$, and highlights that the error is of the order of $10^{-3}$ pointwise. Finally, as in the previous test case, Figure 10 illustrates the loss profiles. Again, the L‑BFGS optimizer converges more quickly and achieves lower loss values than the Adam optimizer.
@@
Figure 8. Transport of a constant density from an ellipse domain to a rotated ellipse domain. Visualization of the transport map $\nabla u_{NN}$.
Figure 9. Transport of a constant density from an ellipse domain to a rotated ellipse domain. Top left: first $x$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,x}-\nabla u_{ex,x}\right|$; Bottom left: second $y$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,y}-\nabla u_{ex,y}\right|$.
Figure 10. Transport of a constant density from an ellipse domain to a rotated ellipse domain. Loss function vs epochs. For the L-BFGS optimizer, the total epoch count is taken as the number of (outer) L‑BFGS iterations times the number of inner sub‐iterations. Shaded bands show the 5th–95th percentile range over ten independent runs.
@@ 4.5. Transport of a Gaussian density in the unit square to another Gaussian density in the unit square
where $c_{0}$ is a normalization constant, $\sigma^{2}=0.25$ and ${x}_{0}=(0.25,0.75)$, and g(x)g({x}) defined similarly, with $\sigma^{2}=0.25$ and ${x}_{0}=(0.75,0.25)$. The histogram in Figure 13 (left) shows the distribution of points sampled from the source Gaussian distribution (4.1) in the unit square and then transported to the target Gaussian distribution through the map $\nabla u_{NN}$ (right). Even if there is no analytical solution, we observe that the distribution of the transported points very closely resembles a Gaussian distribution.
@@ 5. Conclusions and Perspectives
Since we look for a convex solution of the Monge-Ampère equation, input convex neural networks are suitable to obtain appropriate approximations as they are seeking for solutions in the right subspaces. This advantage comes at the cost of increased training complexity.
@@
The relatively high computational expense currently limits the practical problem size and network scale. Furthermore the variability of the results over the samples show that the error coming from the PINNs approach is rather difficult to control. Thus, future avenues of investigations should therefore focus on reducing the CPU overhead of the method, and controlling the error of the method, for instance with hybrid methods. Additional perspectives also include the further incorporation of several sources and targets, via multimodal Gaussian density functions, for real-life applications such as raw material transportation with several sources, as well as the treatment of obstacles. Finally, the investigation of the Monge-Ampère in larger number of space dimensions could be addressed.